OpenStudy (anonymous):

divide the complex numbers 4+5i/2-3i

OpenStudy (anonymous):

just multiply by the conjugate of the denominator (2+3i)/(2+3i)

OpenStudy (anonymous):

conjugate? lol

OpenStudy (anonymous):

\[\large \frac {4+5i}{2-3i} \times \frac {2+3i}{2+3i}\] can you handle the rest?

OpenStudy (anonymous):

the conjugate of a+bi is a-bi ... and vice/versa

OpenStudy (anonymous):

\[\frac{ 4+5i }{2-3i }\\\\ \frac{ 4+5i }{2-3i }*\frac{ 2+3i }{2+3i}\\\\ \frac{ 8+22i+15(i)^2}{4-(3i)^2 }\\\\ \frac{ -7+22i}{ 4+9 }\\\\ \frac{ -7+22i }{13}\\\\ \frac{ -7}{13}+ \frac{ 22i }{13} \]